OpenStudy (anonymous):

write the missing digit to make the number divisable by three 3_,817

OpenStudy (anonymous):

If a number is divisible by 3, its digits add up to a divisor of 3. \[3+8+1+7 = 19\] You will now notice that there is more than one answer. It could either be 2, 5, or 8 because \[3+2+8+1+7 = 21\]\[21/3=7\] \[3+5+8+1+7 = 24\]\[24/3=8\] \[3+8+8+1+7 = 27\]\[27/3=9\]

OpenStudy (anonymous):

thanks